OpenStudy (anonymous):

Which of the following is a factor of f(x) = 5x3 + 24x2 − 75x + 14?

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

i tried grouping it didnt work

OpenStudy (welshfella):

yea - i cant see grouping working here

OpenStudy (anonymous):

so what would work?

OpenStudy (welshfella):

we could try the factor theorem like if f(1) = 0 then x - 1 is a factor but that can be a bit tediuos

OpenStudy (anonymous):

x-1 or x+1?

OpenStudy (welshfella):

try -1 and 1 but thats no good have you got a list of options?

OpenStudy (anonymous):

(x + 1 over 2) (x − 1 over 2) (x + 2) (x − 2)

OpenStudy (welshfella):

f(2) = 0 so that makes (x - 20 a factor

OpenStudy (welshfella):

* (x - 2)

OpenStudy (lynfran):

we can factorize here (x+7)(x-2)(5x-1)

OpenStudy (welshfella):

5(8) + 24(2) - 150 + 14 = 0

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

i have another question

OpenStudy (welshfella):

you're welcome

OpenStudy (anonymous):

will u answer it?

OpenStudy (welshfella):

i'll try

OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 15x2 + 68x − 96. Include the y-intercept, x-intercepts, and the end behavior.

OpenStudy (welshfella):

the y-intercept is easy its value of y when x = 0 = 0 - 0 + 0 - 96

OpenStudy (anonymous):

so its (0,96)?

OpenStudy (welshfella):

you need to factor it to find the x - intercepts

OpenStudy (welshfella):

(0 , -96)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and for the x-imtercept?

OpenStudy (welshfella):

we can try factoring by look at factors of 96

OpenStudy (lynfran):

factor out the function we get (x-3)(x-4)(x-8) now let each bracket equal 0 and solve for x

OpenStudy (anonymous):

so x = 3,4 and 8?

OpenStudy (welshfella):

2*2*2*2*2*3 (x -3) might be a factor f(3) = 0 so that one - oh LynFran has done it already

OpenStudy (anonymous):

so which one of the 3 is my answer?

OpenStudy (lynfran):

all 3 are the x intercept

OpenStudy (anonymous):

oh

OpenStudy (welshfella):

@LynFran = please explain how you got that . Openstudy advises us to guide the poster to the answer

OpenStudy (anonymous):

and what is the end point,

OpenStudy (anonymous):

how do we find the end point?

OpenStudy (anonymous):

how do we find the end point?

OpenStudy (welshfella):

lets look at the highest root which is 8 the graph passes through this point (8,0) so we need to look at what hapenns to f(x) for x > 8 find f(9) and f(10) f(9) = 30 f(10) = 84 - the graph is rising towards infinity at the endpoint

OpenStudy (anonymous):

so the answer is infinity?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ok thanks man

OpenStudy (welshfella):

I was wondering how LynFran was factoring the function so quickly - she was using software to do it for her! Thats no way to learn math - you cant use software in an examination.

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

lol @welshfella